OpenStudy (anonymous):

What value of c makes the trinomial below a perfect square? 4x2 + 16x + c

OpenStudy (anonymous):

4x^2+16x+c (2x)^2 + 2*2x*4 + (4)^2 = (2x+4)^2 = a perfect square so, (a+b)^2 = a^2+2ab+b^2) in our case: a = 2x b = 4 c = 4^2 = 16